Understanding Relationships, Incomplete Data, Overhead Analysis

Lynwood Company produces surge protectors. To help control costs, Lynwood employs a standard costing system and uses a flexible budget to predict overhead costs at various levels of activity. For the most recent year, Lynwood used a standard overhead rate of $18 per direct labor hour. The rate was computed using practical activity. Budgeted overhead costs are $396,000 for 18,000 direct labor hours and $540,000 for 30,000 direct labor hours. During the past year, Lynwood generated the following data: (a) Actual production: 100,000 units; (b) Fixed overhead volume variance: $20,000 U; (c) Variable overhead efficiency variance: $18,000 F; (d) Actual fixed overhead costs: $200,000; and (e) Actual variable overhead costs: $310,000.

Required:

1. Calculate the fixed overhead rate.

2. Determine the fixed overhead spending variance. Enter amount as a positive number and select Favorable or Unfavorable.

3. Determine the variable overhead spending variance. Round hours in the interim calculations to a whole hour. Enter amount as a positive number and select Favorable or Unfavorable.

4. Determine the standard hours allowed per unit of product. Assume actual units equal planned units. Round to five decimal places.
Answer #1

Solution:

1.Calculation of the fixed overhead rate.

Fixed overhead rate = Total overhead rate  - Variable ovrhead rate,

here ,

Total overhead rate = $18 per hour( Given)

Variable overhead rate (by high low method) = Change in overhead cost/ change in labor hours

= ($540,000 - $396,000) / (30,000 hours - 18,000 hours)

=$144,000 / 12,000 hours

=$12 per hour

Therefore,

Fixed overhead rate = $ 18 - $12

= $6 per Hour

b) Calculation of Fixed overhead spending variance.

Fixed overhead spending variance.= Actual Fixed overhead cost - Budgeted Fixed overhead cost

Here

Actual Fixed overhead cost = $200,000 (Given)

Budgeted Fixed Overhead cost = 540,000 - $12 * 30,000

= 540,000 - 360,000

=$180,000

Therefore,

Fixed overhead spending variance. = 200,000 - 180,000

= 20,000 (U)

3. Calcuation of variable overhead spending variance.

Variabe overhead spending Variance = Actual variable overhead - (Variable overhead rate * actual Hours)

Here,

first we need to calculate the Standard hours to calculate Actual hours

Standard Hours = (Budgeted Fixed Overhead -Fixed Ovrhead Volume Variance)/Fixed Overhead rate

= (180,000 - 20,000)/6

=26,667 Hours

Actual Hours = {(Standard Hours * Standard Rate ) - Variable Overhead Efficiency Variance} / Standard Rate

=]{26,667 Hours * $12}- 18000(Given)]/12

=(320,004 -18000)/12

=25,167 Hours

Therfore,

Variabe overhead spending Variance = 310,000 - ($12 * 25,167 Hours)

= 310,000 - 302004

=7996 (u)

4)standard hours allowed per unit of product

Standard Hours allowed per unit = Standard Hours / No of actual units

= 26,667 Hours / 100,000 units

= 0.26667 Hpurs per unit.